Mother age is 3time that of daughter age after 4cyear their combine age is 60 what r their present age?

Answer by addingup(3677)

M= 3d
M+4+d+4= 60 Since the problem says M= 3d, substitute:
3d+4+d+4= 60
4d+8= 60
4d= 52
d= 13 is the age of the daughter
M= 13*3= 39 is the age of the mother